OpenStudy (rissa):

Write y = 5/6x + 1 in standard form using integers.

Directrix (directrix):

5x - 6y = -6 because ..

Directrix (directrix):

y = (5/6) x + 1 6y = 5x +6 -5x + 6y = 6 5x - 6y = -6

Directrix (directrix):

The Standard Form for a linear equation in two variables, x and y, is usually given as Ax + By = C where, if at all possible, A, B, and C are integers, and A is non-negative, and, A, B, and C have no common factors other than 1.

OpenStudy (rissa):

my choices are A. –5x + 6y = 1 B. –5x – 6y = 6 C. –5x + 6y = 6 D. 6x – 5y = 6 though :/

Directrix (directrix):

C. –5x + 6y = 6 is equivalent. My answer multiplied through by -1 will give C. 5x - 6y = -6 --by -1---> -5x + 6y = 6